OpenStudy (anonymous):

FInd the angle in radians between vectors u= cos ((7pi)/6)i +sin ((7pi)/6)j, and v= cos((5pi)/4)i+sin((5pi)/4)j

OpenStudy (anonymous):

For angles \(\vec{a}\) and \(\vec{b}\), the angle \(\theta\) between them can be found using the dot product relation, \[\vec{a}\bullet \vec{b}=\|\vec{a}\|\|\vec{b}\|\cos\theta\]

OpenStudy (loser66):

@SithsAndGiggles Is it not that we just subtract the angles? \(\dfrac{5\pi}{4}-\dfrac{7\pi}{6}=\dfrac{\pi}{12}\)?

OpenStudy (loser66):



OpenStudy (anonymous):

That's another way, yes. We end up with the same result.

OpenStudy (loser66):